OpenStudy (anonymous):

Solve for x: |x + 2| + 16 = 14 x = −32 and x = −4 x = −4 and x = 0 x = 0 and x = 28 No solution

OpenStudy (mathteacher1729):

What is the very first step you did to solve this?

OpenStudy (anonymous):

You would plugin the values they give for x to see if one works

OpenStudy (mathteacher1729):

:: sigh :: I despise multiple choice questions. :p Yes, that is one way of going about it. But if you wish to understand this problem, it would be helpful to graph it and then go through the algebraic steps to understand why the solution works.

OpenStudy (anonymous):

ok

OpenStudy (mathteacher1729):

go to https://www.desmos.com/calculator and type these two equations, then zoom out until you see the functions. y = abs(x+2) + 16 y = 14

OpenStudy (anonymous):

I just did

OpenStudy (mathteacher1729):

Do the lines intersect?

OpenStudy (anonymous):

yes

OpenStudy (mathteacher1729):

To solve abs(x+2) + 16 = 14 subtract 16 from both sides abs(x+2) = -2 NO SOLUTION! The absolute value of ANYTHING is always 0 or greater. So the absolute value of (x+2) will always be 0 or greater. There's NO WAY for it to be -2. So that's that. :) For fun, try typing in y = a*abs(x+b) + c and then press enter, then click "all".

OpenStudy (anonymous):

i did

OpenStudy (mathteacher1729):

Click the sliders around to see what a,b, and c do to the graph. :)

OpenStudy (anonymous):

i did

OpenStudy (mathteacher1729):

You can do that with any graph. Like y = a*x^2+b*x+c y = a*sin(b*x +c) + d y = a*sqrt(b*x+c) + d etc.

OpenStudy (anonymous):

oh ok